OpenStudy (anonymous):

#7 5log(baseB) Q +2log (baseB)Y A.log (baseB)(Q^5Y^2) B.(5+2)log (baseB)(Q+Y) C. Log (base B)(Q^5+Y^2) D.Log (baseB) (QY^5+2)

OpenStudy (anonymous):

well let me try this xD

OpenStudy (anonymous):

solution is A

OpenStudy (anonymous):

\[\log _{b}(p ^{n})=nlog _{b}p\]

OpenStudy (anonymous):

5 log_B Q + 2 log_B Y = log_B Q^5 +log_B Y^2 since log_x Z + log_x Y = log_x ZY then the above can be simplified as ... log_B ((Q^5) * (Y^2)) which is A

OpenStudy (anonymous):

I concur with bnb